OpenStudy (anonymous):

solve the equation given below 18 - 8x - 15 = -7 - 5 ( x - 1) + 5

OpenStudy (anonymous):

i got the answers -5x = 7

OpenStudy (anonymous):

but my answer choices are only A: 6/13 B; 1/2 C: 0 d: 2

Parth (parthkohli):

\[18 - x - 15 = -7 - 5(x - 1) + 5 \]\[\implies 3 - x = -7-5x + 5 + 5 \]\[ \implies 3 - x = -5x + 3\]

Parth (parthkohli):

No, I don't think you are right @mumina1999

Parth (parthkohli):

Can you show your steps?

OpenStudy (anonymous):

ooo i got it so it would be 2 = 2

Parth (parthkohli):

No! You have to solve for \(x\)!

OpenStudy (anonymous):

2x = -2x

OpenStudy (anonymous):

how do i solve for x